Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎


Crew 98 Log Book for January 16, 2011

Photos of the day

Commander's Check-In Report

Virgiliu Pop Reporting

Crew Physical Status: Excellent

Time Departed/Returned from EVA: No official EVA performed today, as we are not yet in sim, but three ATV training trips were performed to Heidi's Hill and back.

Brief Narrative of Field Mission Results: All crew members are now able to drive the ATVs!

EVA Data/Interpretations: N/A

Engineering/Hab Maintenance: As per the engineering report.

Report Transmission Schedule: Intended by 8 PM.

Plans for Tomorrow: Further training, and hab maintenance.

Inventory: N/A

Support Requested: Toilet needs to be fixed.

Miscellaneous: Crews need to be informed as to the location of the ATV registrations, as they were asked for by a ranger.

Commander's Report

Virgiliu Pop Reporting

First morning at the hab, and what a joy! We decided that today we can sleep in, but everybody was up by 9. We snacked for breakfast, then three of us headed to Hanksville, as we are not yet in sim. There we met by chance Judah Epstein, the commander of the previous crew, and we reciprocated his hospitality by inviting him for lunch at the hab. I cooked pasta with tomatoes and mushrooms (normal cooking) and the last bits of Kung-Fu chicken (dehydrated). After lunch, Judah trained Iulia in using the GPS.

Once Judah left, we started the real ATV training - we had three parties (Dragos and Haritina / Dragos, Iulia and Betty / Dragos, Haritina and Virgil), all going to Heidi's Hill and back. We are now fully knowledgeable ATV drivers. The last party was stopped by a friendly ranger who asked to see the registration documents of the vehicles - the matter was referred to mission supoprt, with the request that future crews be informed as to their whereabouts. Meanwhile, Haritina started her recycling programme - we decided that we can recycle the garbage originating at the MDRS and maybe find a helpful local use for some of it, while in the future we can see about composting the food leftovers.

Before dinner, three crew members went to Hanksville and came back in half an hour - at their return they pretended to be wild martian coyotes.

We will soon eat dinner (dehydrated mushroom pilaf, yet another leftover can) and the engineering team will perform their rounds and write the report. Beatrice and Iulia will continue their food inventory.

We have informed mission support that the webcam page has still the details of the previous crew (names and mission patch), while it broadcasts us. The issue has not been solved by the time of this report.

Engineering Report

Haritina Mogosanu Reporting


Kitty (Diesel Generator) (Used? Yes/No): Not used

Honey (Gas Generator) (Used? Yes/No): Yes, running all day

Battery Charge Level:?
28.9 V green
33.8 V armor plate
40.7 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator ran all day ok

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.41

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 1 containers full x 5 gallons, 2 half, 1 a quarter.


Oil: Not checked we are not checking the oil and DG said to leave it on

Kitty Oil Quantity (# of Quarts in storage):

Honey Oil Quantity (# of Quarts in storage):

ATV Oil Quantity (# of Quarts in storage): on the engineering work bench 2 x 946 ml

Car Oil Quantity (# of Quarts in storage):

Water Status:

Outside Potable Water Tank Level (inches from bottom): 12

Trailer Potable Water Tank Level (inches from bottom): less than 1/4 and frozen

Hab Potable Water Tank Level (inches from bottom): 8

Potable Water Meter Reading: not sure where it is, help

In to GreenHab Meter Reading: same



Condition of plants in Tank 1: poor

Condition of plants in Tank 2: poor

Condition of plants in Tank 3: poor

Greenhab Notes/Comments: plants are very few and have some leaves dead.


Telescope Used? (Yes/no) no

Observatory Notes/Comments: Observatory in good condition, dome bolted, ok


Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: not checked
Odometer at end of day: -123134
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: filled
Tire Status: not checked

ATV 2 (Spirit):y
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked

ATV 3 (Viking I):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked

ATV 4 (Viking II):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status:not checked

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 78

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 58

Computers/Networking Infrastructure:

Notes/comments on internet/computers?: all ok

General Engineering Notes/Comments:

Journalist Report

Dragos Bratasanu Reporting

The day began early for some of the crewmembers as I woke at 5 AM in the morning. For others it started at 9 AM. After a quick snack, all the crew members wrote short reports about the first day and press releases for the Romanian media.

The day in the field started at about 11 AM when all members rode the ATVs to get acquainted with the surroundings. The ride was wonderful and the view breathtaking.

The evening got extremely interesting for some of us because we were stopped by the park ranger. I remember when I was a child and I used to watch American movies on TV, especially Yogi Bear. The officer reminded me of the characters in the movies of my youth. However, things turned out all right - he was extremely friendly and visited the hab.

All the team members are writing their evening reports intensely and some of them are trying not to fall asleep while writing. I believe the jet lag is still having fun with them. I will send the pictures of the day shortly.